Needhelpplz:

Guys, I need not the answer but how to do this problem, a.k.a the steps. Also, i need to know why. Alright. Susan begins counting backward from 1298 by 4, saying one number every 5 seconds. At the same time, Jim begins counting forward from 171 by 3, saying 1 number every 45 seconds. What number will they both say at the same time? Please explain thoroughly as this is one of the question for the hunter college exam and going to hunter is my life dream....

Needhelpplz:

It's not an essay

Zepdrix:

lol thinking, chillax :D

lowkey:

this is arithmetic sequence

Needhelpplz:

Can anybody help me?

Needhelpplz:

Please?

Zepdrix:

So like... Susan is counting down from 1298. She's subtracting 4 from that number every 5 seconds. So if x represents time, in seconds, then every x/5 she'll subtract 4 from the number. \[\large\rm 1298-4\left(\frac x5\right)\]

Needhelpplz:

Ok...

Zepdrix:

So ummm

Needhelpplz:

Wait I know one thing

Zepdrix:

If we want this problem to make sense, then she shouldn't be subtracting pieces of 4 at say, x=1, 2, 3, 4 and all these values that are not multiples of 5. So maybe we need to take the floor of this x/5. Ehhh maybe that just makes things more confusing tho :U

Needhelpplz:

The equation is "supposedly" 1298-4x=171+3x

Needhelpplz:

But that's where I got completely confused

Zepdrix:

hmm that doesn't look right :O Especially if Jimbo is adding every 45 seconds. That wasn't a typo, right?

Needhelpplz:

Um no

Needhelpplz:

Wait I think the time does not matter because they are counting at same intervals

Needhelpplz:

That's just what I think

Zepdrix:

They are counting at same intervals ... hmm interesting :D sec thinking lol

Needhelpplz:

Ok

Zepdrix:

Hmm ya that equation does work. I'm having trouble wrapping my head around the fact that we can ignore the 5 second and 45 second counting markers tho, so weird :U hmm

Needhelpplz:

45?

Needhelpplz:

Oh shoot

Needhelpplz:

Srry that was a typo

Needhelpplz:

45 is 5

Zepdrix:

LOL you bozo XD dang that's what I was asking about earlier lol

Needhelpplz:

Oops XD

Needhelpplz:

Sorry

Zepdrix:

So let's assume t is measuring time in seconds. And we'll let x be multiples of 5: \(\large \rm x=5t\) Then every new x (every 5 seconds that passes), we'll subtract 4 from Susan's value. \(\large\rm 1298-4x\) So for example, if x=2 (which means 2 groups of 5 seconds have passed), then we've subtracted 2 fours from our starting value. \(\large \rm 1298 - 4(2)\) So we have our equation for Susan's counting. What do you think so far, does any of this make sense? :U

Zepdrix:

Oh maybe Hero can help XD

Hero:

I think I know what the answer is but I can't say the answer out loud.

Needhelpplz:

Hero?

Needhelpplz:

Uh I know the answer from other sites that I checked but the other sites wouldn't tell me how they got it

Needhelpplz:

The answer is 654

Needhelpplz:

But I dont know how you get it

Zepdrix:

The same can be done for Jim's counting, every 5 seconds (every x), we'll add 3. \(\large\rm 171+3x\)

Zepdrix:

This is the equation for Jim's counting.

Zepdrix:

expression*

Zepdrix:

And we want to know what value of x gives us the same number for each of them. So we'll equate these to one another. \(\large\rm 1298-4x=171+3x\)

Needhelpplz:

Alright

Zepdrix:

Then simply solve for x. To isolate x, we'll get the x-stuff on one side, and everything else on the other side. Subtract 171 from each side, \(\large\rm 1298-171-4x=3x\) Add 4x to each side, \(\large\rm 1298-171 = 3x+4x\)

Zepdrix:

\(\large\rm 1127=7x\)

Zepdrix:

Divide by 7, \(\large\rm x=161\)

Needhelpplz:

Wait

Needhelpplz:

Why do you do the first part

Needhelpplz:

Where you apparently isolate x

Zepdrix:

Susan is counting down. After x amount of time, she arrives at some value. Jim is counting up. After the same x amount of time, he arrives at the same value as Susan. So we set the expressions equal to one another because they represent the same resulting value that they ended up with. And x is the time that passed for them to each arrive at this number.

Zepdrix:

Mmm I dunno, I'm struggling to explain this well :p hmm

Needhelpplz:

Hmmmmmmm.......

Zepdrix:

We're trying to solve for x, the time that passes in order for these "pieces" to be equal. The pieces being: 1298-4x and 171+3x

Needhelpplz:

Yeah, ok

Needhelpplz:

Go on

Zepdrix:

So after we apply those basic Algebra steps, we find that x=161. 161 of these 5 second intervals must pass in order for Susan's number to shrink down to be equal with Jim's rising number.

Zepdrix:

From there, you can plug 161 into either counting expression to see which number they ended up at.

Zepdrix:

Susan: \(\large\rm 1298-4(161)=654\)

Zepdrix:

Or you could plug the value into Jim's expression: \(\large\rm 171+3(161)=654\)

Needhelpplz:

Ok..... only part i dont get is the thing where

Needhelpplz:

Subtract 171 and add 4x

Needhelpplz:

Could you please explain but simply?

Needhelpplz:

Because I dont understand...

Zepdrix:

When you arrive at an equation like this:\[\large\rm 1298-4x=171+3x\]Your goal is to try and get it to look like this:\[\large\rm x=stuff\] We're trying to figure out this special value of x which makes the equation hold true. That requires us to get x alone, get all of the "other stuff" away from the x.

Zepdrix:

We have two separate terms containing x, (4x) and (3x) so it's a lil trickier. We'll try to get both x terms onto one side of the equation We do so because then we can combine like-terms afterwards.

Needhelpplz:

Why does 1298-171=7x equate to 1298=7x? Shouldn't it be 1127=7x

Zepdrix:

Correct, it equates to 1127=7x. I think that's what I wrote above as well.

Needhelpplz:

No but why isn't it1127? Why did the 117 just dissapear

Zepdrix:

Lemme do the Algebra steps again real quick, I'll color the steps for clarity.

Needhelpplz:

Ok

Zepdrix:

\[\large\rm 1298-4x=171+3x\] Add 4x to each side,\[\large\rm 1298-4x\color{orangered}{+4x}=171+3x\color{orangered}{+4x}\]We do this so the -4x and +4x will cancel out on the left.\[\large\rm 1298\cancel{-4x\color{orangered}{+4x}}=171+3x\color{orangered}{+4x}\]\[\large\rm 1298=171+3x\color{orangered}{+4x}\]

Zepdrix:

Then we'll subtract 171 from each side, so that only x-stuff remains on the right side of the equation.

Needhelpplz:

Give me a sec to figure this out..

Zepdrix:

\[\large\rm 1298\color{royalblue}{-171}=171\color{royalblue}{-171}+3x\color{orangered}{+4x}\]

Zepdrix:

We're doing this so the 171-171 will disappear from the right side of the equation,\[\large\rm 1298\color{royalblue}{-171}=\cancel{171\color{royalblue}{-171}}+3x\color{orangered}{+4x}\]\[\large\rm 1298\color{royalblue}{-171}=3x\color{orangered}{+4x}\]

Needhelpplz:

Uh but If on the other side of the equation there is 1298-171 then shouldn't the result be 1127=7x?

Zepdrix:

Yes, our next step would be to apply the subtraction on the left side,\[\large\rm 1127=3x\color{orangered}{+4x}\]And add together the like-terms on the right,\[\large\rm 1127=7x\]

Needhelpplz:

Give me a sec

Zepdrix:

Notice we don't quite have it the form of x=stuff yet. There is still some junk attached to our x.

Needhelpplz:

Ohhhhh I see now

Needhelpplz:

Dude

Zepdrix:

\[\large\rm 1127=7x\]The 7 is multiplying the x, so to get it off the x, we apply the inverse of multiplying. We divide 7.\[\large\rm \frac{1127}{7}=\frac{7x}{7}\]And apply cancellation,\[\large\rm \frac{1127}{7}=\frac{\cancel7x}{\cancel7}\]

Zepdrix:

\[\large\rm 161=x\]

Zepdrix:

One of the weird things to keep in mind, for this problem, ...

Needhelpplz:

?

Zepdrix:

x represents multiples of 5. So if they had asked the question a little differently, say, "How much time must pass for them to get the same value?" Then we would say 161*5 seconds must pass.

Zepdrix:

They didn't ask for that though, so no big deal.

Needhelpplz:

Ok, so then we would do one of the two either 171+161×4 or 1298-161×4 right?

Needhelpplz:

To get 654

Zepdrix:

Correct. Since we expect them to end at the same value, either expression should work.

Zepdrix:

Yayyyy we got it! :D

Needhelpplz:

Thanks man

Needhelpplz:

For spending your time on me

Zepdrix:

np XD

Needhelpplz:

Ok I just became your fan

lowkey:

yes because to become fans you need to be helped